OpenStudy (anonymous):

How many six-letter "words" can be formed from the letters A, B, C, D, E, F, if each letter can be used only once?

Parth (parthkohli):

\(6P6 = 6!\)

Parth (parthkohli):

Another explanation is: You have six slots _ _ _ _ _ You may fill six letters in the first, so six possibilities. Since one letter got taken, then we have five.. it goes down till 1.

Parth (parthkohli):

Now you multiply \(6 \times 5 \times 4 \times 3 \times 2 \times 1\).